if someone wanted an average grade of 90 and they base the grade on 4 tests and the first 3 scores were 84 93 and 88 what would the last test score grade have to be to get a total average of 90?
85
88
92
or
95
: if someone wanted an average grade of 90 and they base the grade on 4 tests and the first 3 scores were 84 93 and 88 what would the last test score grade have to be to get a total average of 90?
85
88
92
or
95

Answer by mangopeeler07(442)

Set up an equation for the average: (84+93+88+x)/4=90

Combine like terms:

(265+x)/4=90

Multiply both sides by 4.

265+x=360

Subtract 265 from both sides

x=95

So for the average grade to be 90, the person would have to get a 95.